17% of students in a statistics class receive an A. out of a sample of 30, what is the probability that 10 students receive an A? no more than 5 students receive an A? what is the expected value?
Answer #1

for binomial distribution ; parameter n=30 and p=0.17

a)

probability that 10 students receive an A =30C10(0.17)10(0.83)20 =0.014582

b)

P( no more than 5 students receive an A )=P(X<=5)= =0.597205

expected value =np=30*0.17 =5.1